Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Two Boxes

Expert replies
by sk8ternite » Tue Aug 25, 2009 11:07 am
There are two boxes of markers. The first box has only red, orange and yellow markers. The second box has only yellow, green and blue markers. If a marker is chosen at random from each of the boxes, is the probability that both markers will be yellow less than 30%?
(1) Four of the 12 total markers are yellow
(2) The probability of selecting a red or orange marker is 3/5 and the probability of selecting a green or blue marker is 2/7 .
Join the discussion
Source: — Data Sufficiency |

Re: Two Boxes

by maihuna » Tue Aug 25, 2009 11:15 am
sk8ternite wrote:There are two boxes of markers. The first box has only red, orange and yellow markers. The second box has only yellow, green and blue markers. If a marker is chosen at random from each of the boxes, is the probability that both markers will be yellow less than 30%?
(1) Four of the 12 total markers are yellow
(2) The probability of selecting a red or orange marker is 3/5 and the probability of selecting a green or blue marker is 2/7 .
B: 2/5*5/7 = 2/7 = <30%
Charged up again to beat the beast :)
Join the discussion

Re: Two Boxes

by sk8ternite » Tue Aug 25, 2009 11:53 am
maihuna wrote:
sk8ternite wrote:There are two boxes of markers. The first box has only red, orange and yellow markers. The second box has only yellow, green and blue markers. If a marker is chosen at random from each of the boxes, is the probability that both markers will be yellow less than 30%?
(1) Four of the 12 total markers are yellow
(2) The probability of selecting a red or orange marker is 3/5 and the probability of selecting a green or blue marker is 2/7 .
B: 2/5*5/7 = 2/7 = <30%
https://www.beatthegmat.com/a/2009/08/25 ... n-pitfalls

How come A cant be right, Ive tried different combinations and always get less than 30%. We know that there are 4 yellow markers and 12 total markers, so creating different scenarios, I always get under 30%, unless im doing something wrong?
Join the discussion

by pradeepsarathy » Tue Aug 25, 2009 12:38 pm
IMO D

Stmt 1 -

Probability of selecting yellow markers can be found - 4C2 / 12C2
Sufficient

Stmt 2 -

Probability of selecting a yellow marker from the first box is 1 - 3/5 = 2/5
Probability of selecting a yellow marker from the second box is 1 - 2/7 = 5/7

Hence probabiltiy of selecting yellow markers from both the box is 2/5 * 5/7 = 2/7
Sufficient
Join the discussion